COUNTING INVOLUTIONS

Michael Aschbacher, Ulrich Meierfrankenfeld, and Bernd Stellmacher

California Institute of Technology Michigan State University Christian-Albrechts-Universität zu Kiel

There are a number of results on finite groups that are proved by counting involutions; the Brauer-Fowler Theorem [BF] and the Thompson Order Formula (cf. 45.6 in [FGT]) are perhaps the two most important examples. In a less well known paper [B], Bender introduces an involution counting technique useful in analyzing small groups where more traditional local analysis is often ineffective. In particular Bender’s approach can sometimes be used to calculate the order of a group with one class of involutions; we recall that the Thompson Order Formula does not apply to such groups.

In this note we use Bender’s approach to prove:

Theorem 1. Assume M is a nontrivial TI-subgroup of the finite group G, let M^{∗} = N_{G}(M), and assume M^{∗} = MC_{M^{∗}}(z) for some involution z ∈ M^{∗} and C_{G}(x) is of odd order for each x ∈ M^{#}. Then one of the following holds:

(1) M ⊴ G.

(2) n(G) ≤ |G : M|, where n(G) denotes the number of involutions in G.

(3) G ≅ P GL_{2}(m) or L_{2}(m) for some power m of some prime p, M is an elementary abelian p-group of order m, and m ≡ 1 mod 4 if G ≅ L_{2}(m).

(4) G ≅ L_{2}(2^{e}) for some integer e > 1 and M is cyclic of order 2^{e} + 1.

(5) There exists a positive integer d such that G is the split extension of an elementary abelian group of order 2^{2d} by M^{∗}, and M^{∗} is dihedral of order 2(2^{d} + 1).

The first author’s work was partially supported by NSF-9901367

2 MICHAEL ASCHBACHER, ULRICH MEIERFRANKENFELD, AND BERND STELLMACHER

See [FGT] for the definition of basic notation and terminology. For example T is a TI-subgroup of G if distinct conjugates of T in G intersect trivially; O(G) is the largest normal subgroup of G of odd order; G^{#} = G − {1} is the set of nontrivial elements of G.

Notice that if conclusion (2) holds then |M| ≤ |C_{G}(t)| for each involution t in G, while if G has one class of involutions then conclusion (2) of Theorem 1 is equivalent to |M| ≤ |C_{G}(z)|.

The next lemma provides subgroups to which Theorem 1 can be applied. It is a slight variation on the penultimate theorem in [Br]. Given a group G, define the commuting graph of G to be the graph Δ with vertex set G^{#} and g adjacent to h if gh = hg. Write d(g, h) for the distance from g to h in Δ and let Δ^{i}(g) = {h ∈ Δ : d(g, h) = i}.

Lemma 2. Let G be a finite group, z an involution in G, and 1 ≠ x an element of G inverted by z. Assume

(i) d(z, x) ≥ 3, and

(ii) if d(z, x) = 3 then z inverts no member of Δ^{2}(z) ∩ Δ(x). Set M = C_{G}(x). Then

(1) M = C_{G}(y) for each y ∈ M^{#}, so d(z, x) = ∞.

(2) M is a TI-subgroup of G.

(3) M is abelian of odd order and inverted by z.

(4) M^{∗} = N_{G}(M) is a Frobenius group with kernel M and complement C_{M^{∗}}(z).

(5) M is a Hall subgroup of G.

In a group G with one class of involutions, if one has good control over the centralizer of an involution then one can usually obtain strong information about the centralizers of elements at distance 1 from involutions, and hence also some information about elements at distance 2. Lemma 2 and Theorem 1 supply information about centralizers of elements at distance greater than 2 that are inverted by involutions.

Finally to illustrate how Theorem 1 and Lemma 2 can be applied, we give a short, elementary proof of the following well known result:

Theorem 3. Let G be a finite group containing an involution z such that T = C_{G}(z) is dihedral of order 8. Then one of the following holds:

(1) G = T O(G).

(2) G ≅ S_{4} ≅ P GL_{2}(3).

(3) G ≅ S_{5} ≅ P GL_{2}(5).

(4) G ≅ L_{3}(2) ≅ P SL_{2}(7).

(5) G ≅ A_{6} ≅ P SL_{2}(9).

Bender gives a similar proof of a slightly weaker result in [B]. Theorem 1 makes possible only a small simplification in Bender’s treatment, but such simplifications become more crucial in larger examples.

Local characterizations like Theorem 2 of small groups in the existing literature typically involve extensive use of exceptional character theory and block theory. In addition to supplying simplified proofs for such results, another advantage to an approach involving involution counting is the elimination or reduction in appeals to character theory.

Section 1. A lemma of Bender

Let G be a finite group and J the set of involutions in G. For S ⊆ G, let n(S) = |J ∩S| be the number of involutions in S. Following Bender in [B], given a subgroup M of G, define

f = f(G, M) = \frac{n(G)}{|G : M|} − 1.

Observe

(1.1) n(G) > |G : M| iff f(G, M) > 0.

Write G/M for the coset space of cosets Mg, g ∈ G, and represent G by right multiplication on G/M.

Again following Bender, given a nonnegative integer m, define

b_{m} = b_{m}(G, M) = |{C ∈ G/M − {M} : n(C) = m}|.

Observe that for each m, M acts on the set of cosets C of M with n(C) = m.

The following lemma of Bender in [B] is one of the fundamental tools in this paper; its proof is easy and elementary.

(1.2) (Bender) Assume M ≤ G with n(G) > |G : M|. Then

b_{1} = f^{−1}(n(M) + ∑_{i>1}(i − 1)b_{i} − 1 − b_{0}) − 1 − b_{0} − ∑_{i>1} b_{i}.

Section 2. The proof of Theorem 1

Throughout this section we assume the hypotheses of Theorem 1. Continue the notation of the previous section. In addition for x ∈ G let I(x) be the set of involutions in G inverting x. Set m = |M|. Recall m_{2}(G) is the 2-rank of G; that is m_{2}(G) = k where 2^{k} is the maximal order of an elementary abelian 2-subgroup of G.

(2.1) (1) M is abelian of odd order and inverted by z.

(2) I(y) ⊆ Mz = z^{M} for all y ∈ M^{#}.

(3) J ∩ M^{∗} = z^{M}.

(4) m_{2}(M^{∗}) = 1.

(5) C_{M^{∗}}(z) is a complement to M in M^{∗} and [z, M^{∗}] = M. Proof. Let W = C_{G}(M). By hypothesis C_{G}(y) is of odd order for each y ∈ M^{#}, so m is odd, W is of odd order, and C_{M}(i) = 1 for each involution i ∈ M^{∗}. Thus (1) holds. If i ∈ I(y) then i ∈ M^{∗} as M is TI. Then i inverts M, so zi ∈ W. Hence as |W | is odd, i ∈ z^{W}. But by hypothesis, M^{∗} = MC_{M^{∗}}(z), so W = MC_{W} (z), and hence i ∈ z^{M} = Mz. Therefore (2) holds. Further as C_{M}(i) = 1 for each i ∈ J ∩ M^{∗}, (2) implies (3). As m_{2}(M⟨z⟩) = 1, (3) implies (4). As M^{∗} = MC_{M^{∗}}(z) and C_{M}(z) = 1, (5) holds.

(2.2) (1) Mz is the unique coset C of M such thatn(C) > 1.

(2) n(Mz) = m. Proof. Suppose C ∈ G/M with n(C) > 1. Then there are distinct involutions i, j ∈ C. Then ij ∈ M^{#} is inverted by i and j, so by 2.1.2, i, j ∈ I(ij) ⊆ Mz; that is C = Mz, so (1) holds. Part (2) follows from parts (1) and (2) of 2.1.

During the remainder of the section we assume that neither conclusion (1) nor conclusion (2) of Theorem 1 hold. Thus n(G) > |G : M| and G ≠ M^{∗} = N_{G}(M). Let a = |M^{∗} : M|, r = |C_{G}(z)|/a and N = |G : M^{∗}|. Observe a = |C_{M^{∗}}(z)|.

Recall the definition of the parameters b_{i} = b_{i}(G, M) from the previous section.

(2.3) (1) b_{1} = f^{−1}(m − (b_{0} + 2)) − (b_{0} + 2), and

(2) b_{0} < m − 2. Proof. As n(G) > |G : M|, we can apply 1.1 to conclude f > 0 and appeal to 1.2. By 2.2, b_{i} = 0 except when i is 0, 1, or m; also n(M) = 0 and b_{m} = 1. Thus (1) follows from 1.2. As b_{1} ≥ 0, (1) implies (2).

(2.4) Let g ∈ G − M^{∗} and set D = M^{∗} ∩ M^{∗g}. Then (1) n(Mg) = 1 and n(M^{∗}g) = a.

(2) D is a complement to M in M^{∗} inverted by the involution t ∈ Mg, D contains a unique involution i, and D = C_{M^{∗}}(i).

(3) M ∩ M^{∗g} = 1.

(4) N − 1 = m(r − 1).

(5) |G| = Nma = (m(r − 1) + 1)ma.

(6) b_{1} = am(r − 1).

Proof. If n(Mg) ≠ 1 then n(Mg) = 0 by 2.2.1. However as M ∩ M^{g} = 1, the orbit on G/M of Mg under M is of length m, so b_{0} ≥ m, contrary to 2.3.2. Further M^{∗}g is the union of a cosets of M, so n(M^{∗}g) = a, establishing (1).

Next by 2.1.3, z^{G} ∩ M^{∗} = z^{M^{∗}}, so C_{G}(z) is transitive on the set Γ of fixed points of z on G/M^{∗}. (cf. 5.21 in [FGT]) Thus |Γ| = |C_{G}(z) : C_{M^{∗}}(z)| = |C_{G}(z)|/a = r. Let U = M^{∗}g ∩J; by (1), U is of order a. Thus for j ∈ U, the set U of elements of D inverted by j is {uj : u ∈ U} of order a. By (1), U is a set of coset representatives for M in M^{∗}g, so as M^{∗}j = M^{∗}g and D ≤ M^{∗}, we conclude:

M^{∗} = ⋃_{u∈U} Muj = MU = MD.

In particular D is of even order so as j acts on D, D contains an involution i centralizing j. Now by 2.1.5, i inverts M, so M ∩ D = [M ∩ D, i]. Also by 2.1.5 applied to M^{∗g}, [M ∩ D, i] ≤ [M^{∗g}, i] ≤ M^{g}, so M ∩ D ≤ M ∩ M^{g} = 1. Thus D is a complement to M in M^{∗}, so |D| = a = |U| and hence D = U. As D = U, D is inverted by j, so D is abelian. Thus D ≤ C_{M^{∗}}(i), so D = C_{M^{∗}}(i) and as m_{2}(M^{∗}) = 1, i is the unique involution in D. Therefore (2) and (3) hold.

For each g ∈ G − M^{∗}, M^{∗} ∩ M^{∗g} is the stabilizer in M^{∗} of M^{∗}g ∈ G/M^{∗}, so by (2), M^{∗}g is fixed by a unique involution of M^{∗}. Thus by 2.1.3, G/M^{∗} − {M^{∗}} is partitioned by the m sets F ix_{G/M^{∗}}(z^{x}) − {M^{∗}}, x ∈ M. This establishes (4).

Next |G : M^{∗}| = N and |M^{∗}| = ma, so (5) follows from (4). Similarly |G : M| = aN, while by (1), |G : M| = a + b_{1}. Then (4) implies (6).

(2.5) (1) Either (i) For some integer e > 2, C_{G}(z) ≅ E_{2^{e}}, a = 2, and r = 2^{e−1}, or (ii) r = 2 and C_{M^{∗}}(z) is inverted by an involution in C_{G}(z) − C_{M^{∗}}(z).

(2) In case (i):

(a) |G| = 2m(m(2^{e−1} − 1) + 1).

(b) |G : C_{G}(z)| = m(m(2^{e−1} − 1) + 1)/2^{e−1}.

(c) |J | = m(2^{e} − 1).

(3) In case (ii):

(a) |G| = m(m + 1)a.

(b) |G : C_{G}(z)| = m(m + 1)/2.

(c) |J | = m(a + 1).

Proof. If C_{G}(z) ≤ M^{∗} then r = 1, so N = 1 by 2.4.4, contradicting G ≠ M^{∗}. Thus there is g ∈ C_{G}(z) − M^{∗}. Let A = M^{∗} ∩ M^{∗g}; then C_{M^{∗}}(z) = A by 2.4.2. By 2.4.1, there is an involution t ∈ Mg, and by 2.4.2, t inverts A. Further g ∈ C_{M}(z)t = {t}, so each element of C_{G}(z) − A is an involution inverting A. Thus (1ii) holds if r = 2. If r > 2 there exists h ∈ C_{G}(z) − A with hg ∉ A. Thus hg inverts and centralizes A, so A is of exponent 2, and then A = ⟨z⟩ by 2.1.4, so (1i) holds. Hence (1) is established.

Observe that as |C_{G}(z)| = ar, |G : C_{G}(z)| = Nm/r by 2.4.5. Also by 2.4.1 and 2.4.6, |J | = n(M) + b_{1} = m + am(r − 1). Then (2) and (3) are easy calculations, given 2.4.

(2.6) Assume case (ii) of 2.5.1 holds and let A = C_{M^{∗}}(z). Then (1) G is 2-transitive on G/M^{∗} and M is regular on G/M^{∗} − {M^{∗}}.

(2) A is cyclic and semiregular on M.

(3) a = (m − 1)/2 or m − 1.

(4) m is a power of a prime p and M is an elementary abelian p-group.

(5) If a = m − 1 then G has two classes of involutions and G ≅ P GL_{2}(m).

(6) If a = (m − 1)/2 then G has one class of involutions and G ≅ L_{2}(m) with m ≡ 1 mod 4.

Proof. By 2.5.3, |G : M^{∗}| = m + 1, while by 2.4.3, M is semiregular on G/M^{∗} − {M^{∗}}, so (1) holds. Suppose 1 ≠ X ≤ A with 1 ≠ C_{M}(X). As C_{M}(z) = 1, z ∉ X, so as m_{2}(A) = 1, X is of odd order. By (1), C_{M}(X) is regular on F ix(X)−{M^{∗}} and by 2.5.2 there is an involution t ∈ G−M^{∗} inverting X, so C_{M^{t}}(X) is regular on F ix(X)−{M^{∗}t}. Thus Y = ⟨C_{M}(X), C_{M^{t}}(X)⟩ is 2-transitive on F ix(X) and Y ≤ C_{G}(X). Thus there is g ∈ C_{G}(X) with cycle (M^{∗}, M^{∗}t). This is impossible as A⟨t⟩ is the global stabilizer of this set and each g ∈ At inverts X, so g ∉ C_{G}(X) as X is of odd order. Thus A is semiregular on M, so as A is abelian, A is cyclic; that is (2) holds.

By 2.5.3, |J | = m(a+1) and |z^{G}| = m(m+1)/2, so m(m+1)/2 ≤ m(a+1), and hence a ≥ (m − 1)/2. But by (2), a divides m − 1, so (3) holds. Further if a = (m − 1)/2 then J = z^{G}, so G has one class of involutions and m ≡ 1 mod 4 as a is even. If a = m − 1, then A is regular on M^{#}, so G is sharply 3-transitive on G/M^{∗} by (1). In any event, A is irreducible on M, so (4) holds.

In each case G is determined up to isomorphism by a result of Zassenhaus [Z], but we do not need Zassenhaus’ result. Rather we argue as follows:

Suppose a = m − 1. Regard M as an s-dimensional vector space over F_{p}; then A is a cyclic subgroup of GL(M) regular on M^{#}, so A is determined up to conjugation in GL(M). Hence M^{∗} = MA and its action on Ω = G/M^{∗} are determined up to equivalence. Let S = Sym(Ω). In particular a generator x of A is an a-cycle on Ω−{M^{∗}}, so C_{S}(A) = A × ⟨τ⟩, where τ is a transposition in S with cycle (M^{∗}M^{∗}t), and N_{S}(A) = ⟨τ⟩×B, where B = N_{S}(A)_{M^{∗}} is the split extension of A by Aut(A). Thus A is of index 2 in a unique dihedral subgroup D of S such that D ≰ B. Therefore G = ⟨M^{∗}, t⟩ = ⟨M^{∗}, D⟩ is determined up to conjugation in S. Then as P GL_{2}(m) satisfies the hypotheses of G, it follows that G ≅ P GL_{2}(m), so (5) holds.

So assume a = (m − 1)/2. Then G has one class of involutions and satisfies the hypotheses of Theorem 3.5 in Chapter 13 of [G]. Then that result says G ≅ L_{2}(m), so (6) holds.

(2.7) Assume case (i) of 2.5.1 holds. Then (1) Either

(a) G has one class of involutions, m = 2^{e} + 1, and |G| = 2^{e}(2^{2e} − 1), or

(b) G has more than one class of involutions, m = 2^{d} + 1, and |G| = 2^{2d+1}(2^{d} + 1), where d = e − 1.

(2) If G has one class of involutions then G ≅ L_{2}(2^{e}).

(3) If G has more than one class of involutions then G = RM^{∗}, where R = O_{2}(G) ≅ E_{2^{2d}}.

(4) M is cyclic.

Proof. By 2.5.2:

\frac{m(m(2^{d} − 1) + 1)}{2^{d}} = |z^{G}| ≤ |J | = m(2^{d+1} − 1),

so

m(2^{d} − 1) ≤ 2^{d}(2^{d+1} − 1) − 1 = (2^{d} − 1)(2^{d+1} + 1),

and hence

(*)  m ≤ 2^{d+1} + 1 = 2^{e} + 1.

Indeed if G has one class of involutions then the inequality in (*) is an equality, so m = 2^{e} + 1 and hence |G| = 2^{e}(2^{2e} − 1) by 2.5.2, so (1a) holds in this case. Further C_{G}(z) ≅ E_{2^{e}} in 2.5.1.i, so as G has one class of involutions, C_{G}(i) ≅ E_{2^{e}} for each i ∈ J. Therefore (2) holds by Exercise 16.1 in [FGT]. In particular the unique subgroup of G of order 2^{e} + 1 is cyclic, so (4) holds.

Thus we may assume G has more than one class of involutions. Then the inequality in (*) is strick; that is m < 2^{e} + 1. However by 2.5.2,

|G : C_{G}(z)| = \frac{m(m(2^{d} − 1) + 1)}{2^{d}}

so

0 ≡ m(2^{d} − 1) + 1 ≡ 1 − m mod 2^{d},

and hence m = 2^{d} + 1. Thus (1b) holds by 2.5.2.

Let i ∈ J − z^{G}; then i inverts no element of M^{#} as M is a TI-subgroup of G and all involutions in M^{∗} are in z^{G}. Indeed as |G| = (2^{d} + 1)2^{2d+1} and M is a TI-subgroup, all elements in G^{#} of odd order are in conjugates of M, so i inverts no element of G^{#} of odd order. Let R = O_{2}(G); it follows from the Baer-Suzuki Theorem (cf. 39.6 in [FGT]) that i ∈ R. As C_{G}(x) is of odd order for each x ∈ M^{#}, M is semiregular on R. Thus if V is a nontrivial M-chief section on R, |V^{#}| ≡ 0 mod 2^{d} + 1, so |V | ≥ 2^{2d}. Then as |G| = 2^{2d}|M^{∗}|, it follows that R ≅ E_{2^{2d}} and M^{∗} is a complement to R in G. Thus (3) holds. Finally as M is abelian and semiregular on R, (4) holds.

Observe that 2.5-2.7 complete the proof of Theorem 1.

Section 3. The proof of Lemma 2

In this section we assume the hypotheses of Lemma 2. By (i), C_{M}(z) = 1, so z inverts M and (3) holds. Let y ∈ M^{#}. Then d(z, y) ≥ d(z, x) − 1, so if d(z, x) > 3 then d(z, y) ≥ 3. On the other hand if d(z, x) = 3 then as z inverts y, d(z, y) ≥ 3 by (ii). As M is abelian, M ≤ C_{G}(y) and as d(z, y) ≥ 3, z inverts C_{G}(y), so C_{G}(y) is abelian. Thus C_{G}(y) ≤ M, so (1) holds. Then (1) implies (2).

Let i be an involution in M^{∗}. By (1), C_{M}(i) = 1, so i inverts M. Thus iz ∈ C_{G}(M) = M, so i ∈ Mz. Thus J ∩ M^{∗} = Mz = z^{M}, so C_{M^{∗}}(z) is a complement to M in M^{∗} by a Frattini argument. Further for r ∈ C_{M^{∗}}(z)^{#}, C_{M}(r) = 1 by (1), so (4) holds. Finally by (4), M is a Hall subgroup of M^{∗}, while for p ∈ π(M) and P ∈ Syl_{p}(M), N_{G}(P) ≤ M^{∗} by (2), so P ∈ Syl_{p}(G). Thus (5) holds.

This completes the proof of Lemma 2.

Section 4. The proof of Theorem 3

In this section we assume the hypotheses of Theorem 3. As ⟨z⟩ = Z(T) is characteristic in T and T ∈ Syl_{2}(C_{G}(z)), T ∈ Syl_{2}(G). Let A_{1} and A_{2} be the two 4-subgroups of T and let G_{i} = N_{G}(A_{i}). If z is strongly closed in T with respect to G, then applying Thompson transfer to the cyclic subgroup of index 2 in T, O^{2}(G) has cyclic Sylow 2-subgroups. Therefore (1) holds. (cf. 39.2 in [FGT]) Thus we may assume z^{g} ∈ A_{1} − ⟨z⟩.

As ⟨z⟩ = [T, A_{1}] and ⟨z^{g}⟩ = [T^{g}, A_{1}], H = ⟨T, T^{g}⟩ ≤ G_{1} and H induces L_{2}(2) on A_{1}. Thus as A_{1} = C_{G}(A_{1}), G_{1} = H ≅ S_{4}. If A_{1} is strongly closed in T with respect to G, then by Thompson transfer, O^{2}(G_{1}) = O^{2}(G) ∩ G_{1}. Let L = O^{2}(G). Then A_{1} ∈ Syl_{2}(L) and L has one class of involutions with C_{L}(z) = A_{1} ≅ E_{4}. It follows from Exercise 16.1 in [FGT] that L ≅ A_{4} or A_{5}, and then that (2) or (3) holds. Thus we may assume A_{1} is not strongly closed in T with respect to G, so G has one class of involutions and by symmetry, G_{2} ≅ S_{4}.

Suppose x ∈ G^{#} is of odd order and inverted by an involution. As G has one class of involutions, we may assume z inverts x. As C_{G}(z) is a 2-group, d(z, y) = ∞ for each y ∈ G^{#} of odd order. Thus we conclude from Lemma 2 that X = C_{G}(x) is a TI-subgroup of odd order inverted by z and satisfies the hypotheses of the group “M” in Theorem 1. Therefore by Theorem 1, either |X| ≤ |C_{G}(z)| = 8, or G satisfies one of conclusions (3)- (5) of Theorem 1. In the latter case as G has one class of involutions and C_{G}(z) ≅ D_{8}, we conclude (4) or (5) holds. Thus we may assume |C_{G}(x)| = 3 or 5 for each x ∈ G^{#} of odd order inverted by an involution. In particular if X_{i} ∈ Syl_{3}(G_{i}) then as X_{i} is inverted by an involution in T, X_{i} = C_{G}(X_{i}), so N_{G}(X_{i}) = N_{G_{i}}(X_{i}) ≅ S_{3}.

Let M = G_{1}; as |M| > |T |, we may apply 1.2 to M. As G has one class of involutions,

f = \frac{|G : T |}{|G : M|} − 1 = \frac{|M|}{|T |} − 1 = 2.

Further if y ∈ M^{#} then either N_{G}(⟨y⟩) ≤ M, or y is one of the six involutions in M −A_{1}. Thus if u is an involution in G − M then n(Mu) > 1 iff u centralizes one of these six involutions y, in which case ⟨y⟩ = M ∩ M^{u}, so n(Mu) = 2. Hence b_{m} = 0 for m > 2 and b_{2} = 6. Then as n(M) = 9, it follows from 1.2 that

b_{1} = f^{−1}(n(M) + b_{2} − b_{0} − 1) − b_{2} − b_{0} − 1 = \frac{9 + 6 − b_{0} − 1}{2} − 6 − b_{0} − 1 = \frac{−3b_{0}}{2} .

Thus as b_{1} ≥ 0, we conclude b_{1} = b_{0} = 0. Therefore |G : M| = 1 + b_{2} = 7. Now we conclude G ≅ L_{3}(2) by any one of a number of means. For example G is 2-transitive on G/G_{i} for i = 1, 2, so the coset geometry of G on the family {G_{1}, G_{2}} is a projective plane of order |G_{i} : G_{1} ∩ G_{2}| − 1 = 2. As it is an elementary exercise to show there is a unique plane of order 2, G = L_{3}(2) is the group of automorphisms of that plane.

This completes the proof of Theorem 3.

References

[B] H. Bender, Finite groups with large subgroups, Illinois J. Math. 18 (1974), 223–228.

[Br] R. Brauer, On the structure of groups of finite order, Proc. Inter. Cong. Math. Amsterdam (1954), 209–217.

[BF] R. Brauer and K. Fowler, Groups of even order, Annals Math. 62 (1955), 565–583.

[FGT] M. Aschbacher, Finite Group Theory, Cambridge Univ. Press, Cambridge, 1986.

[G] D. Gorenstein, Finite Groups, Harper and Row, New York, 1968.

[Z] H. Zassenhaus, Kennzeichnung endlicher linearer Gruppen als Permutationsgruppen, Hamburg Abh. 11 (1936), 17–40.

Pasadena, California 91125

East Lansing, Michigan 48824

24118 Kiel, Germany